MEDICATION CALCULATION #56

Order: 1000 mL of Lactated Ringer's (LR) solution to be administered in 6 hours

Available: 1 L LR Macrodrip set labeled 10 gtt/mL

a. How many mL of LR should be infused in 1 hour?

b. How many drops per minute (gtt/min) should the client receive?

Please show your step by step calculations.
Answer #1

ANSWERS

   a. mL of Ringer lactate administered in 1 hr

                        Formula

                             ml/hr = Total volume / Total time

                         Here

                              Total volume = 1000 ml.

                              Total time     = 6 hr.

                                ml / hr        = 1000 /6

                                                  = 166.66 ml /hr.

b.Drops per minute (gtt /min)

                   Formula

                         Drops / min = ml / hr * drop factor / 60

                   Here

                          ml / hr = 166.66 ml.

                  Drop factor = 10 gtt/ min.

                  so                   = 166.66 * 10 /60

                                         = 166.66 8 * 0.166

                                         = 27.7 gtt /min